1.A vial of isophane insulin for subcutaneous injection contains 100 units/ml. How many units are there in a 0.7 ml dose?

2.The initial adult dose of metformin, an antihyperglycemic drug, is 1000 mg/day. This dose is increased in increments of 500 mg/day each week until therapeutic results are achieved. If this increase occurs over a period of 2 weeks, what will be the dose at the beginning of the third week?

3.Pioglitazone increases the body’s sensitivity to insulin. When used alone, an initial adult dose of pioglitazone can start at 15 mg/day. Suppose that, after one week, the dosage is increased by 15 mg/day, and that after the second week, another increase by 15 mg/day occurs. How many 15 mg tablets would be needed for three weeks?
Answer #1
  1. Vial of isophane insulin contains= 100 units/ml.

How many units are there in a 0.7 ml dose?

1 ml=100 units

0.7 ml= 70 units

2. The initial adult dose of metformin, an antihyperglycemic drug =1000 mg/day. This dose is increased in increments of 500 mg/day each week until therapeutic results are achieved.

Given that the increase occurs over a period of 2 weeks.

Thus the dose at the beginning of the third week will be = 1000mg + 500 mg after first week+ 500 mg after 2nd week (i.e. the beginning of the third week)

= 2000 mg / day.

  1. Initial adult dose of pioglitazone = 15 mg/day.

After one week, the dosage is increased by 15 mg/day,

Then again after the second week, another increase by 15 mg/day occurs.

Then total dose for one week =15X7=105 mg

Then in second week dose will be 15+15=30mg/day

then total dose in second week =30X7=210 mg

Similarly in third week dose will be =30+15=45 mg

Then in third week total dose will be=45X 7=315 mg

Grand total dose in mg of all three weeks =105+210+315= 630 mg

How many 15 mg tablets would be needed for three weeks?

15 mg in 1 tablet

1mg in=1/15 tablets

630 mg will be from=630/15= 42 tablets will be required for whole doses for three weeks.